Logiweb codex of hello world

Up Help

hello world,

hello world

[hello world →^{pyk} “hello world”]

The pyk compiler, version 0.grue.20050502 by Klaus Grue GRD-2005-06-22.UTC:14:45:04.495547 = MJD-53543.TAI:14:45:36.495547 = LGT-4626168336495547e-6